OpenStudy (punderson):

Choose the missing step in the given solution to the inequality -6x - 10 > 14 + 2x.(5 points) -6x - 10 > 14 + 2x -8x > 24 x < -3

rebeccaxhawaii (rebeccaxhawaii):

what are you asking

OpenStudy (punderson):

What step is missing in that problem @rebeccaxhawaii

OpenStudy (punderson):

Here are the choices -4x - 10 > 14 -8x - 10 > 14 -4x - 10 < 14 -8x - 10 < 14

OpenStudy (dogzcatz):

-4x - 10 > 14

OpenStudy (punderson):

Can you help me with more questions @DogzCatz

OpenStudy (dogzcatz):

Sure man

OpenStudy (punderson):

ok heres this one- @DogzCatz The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F. : C = 5 over 9 (F - 32) Which of the following formulas correctly solves for F? Choices- F = 9 over 5 C + 32 F = 9 over 5 C - 32 F = 9C + 32 over 5 F = 9C - 32 over 5

OpenStudy (dogzcatz):

f=9/5c+32

OpenStudy (punderson):

Solve for L: A = L • W • H Choices- @DogzCatz L = A times W over H L = W times H over A L = A over quantity W times H L = W over quantity H times A

OpenStudy (dogzcatz):

L = W times H over A

OpenStudy (punderson):

Few more ok! Solve for x. -ax + 2b > 8 Choices- @DogzCatz x < the quantity negative 2 times b plus 8 all over negative a x > the quantity negative 2 times b plus 8 all over negative a x < the quantity 2 times b minus 8 all over negative a x > the quantity negative 2 times b minus 8 all over a Which of the following is not a way to represent the solution of the inequality 5(x + 2) greater than or greater to 7x + 2(x - 1)? Choices- A number line with a closed circle on 3 and shading to the right A number line with a closed circle on 3 and shading to the left 3 greater than or equal to x x less than or greater to 3 A company did a quality check on all the packs of nuts it manufactured. Each pack of nuts is targeted to weigh 18.25 oz. A pack must weigh within 0.36 oz of the target weight to be accepted. What is the range of rejected masses, x, for the manufactured nuts? Choices- x < 17.89 or x > 18.61 because |x - 18.25| > 0.36 x < 17.89 or x > 18.61 because |x - 0.36| + 18.25 > 0 x < 18.25 or x > 18.61 because |x - 18.25| > 0.36 x < 18.25 or x > 18.61 because |x - 0.36| + 18.25 > 0

OpenStudy (punderson):

The table below shows two equations: Equation 1 |4x - 3|− 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? Choices- Equation 1 and equation 2 have no solutions. Equation 1 has no solution, and equation 2 has solutions x = -4, 1. The solutions to equation 1 are x = 3, -1.5, and equation 2 has no solution. The solutions to equation 1 are x = 3, -1.5, and equation 2 has solutions x = -4, 1.

OpenStudy (dogzcatz):

For -ax + 2b > 8 I got x<2ba−8a

OpenStudy (dogzcatz):

A number line with a closed circle on 3 and shading to the left

OpenStudy (dogzcatz):

x < 17.89 or x > 18.61 because |x – 18.25| > 0.36

OpenStudy (punderson):

For the answers can you do A,B,C,D Please from my order! @DogzCatz

OpenStudy (dogzcatz):

The solutions to equation 1 are x = 3, -1.5, and equation 2 has no solution.

OpenStudy (dogzcatz):

 this is the 2nd equation